OpenStudy (anonymous):

-5√192) -2√243) -8√108) please help subtract terms. Need to show work

OpenStudy (anonymous):

You need to find the least common multiple of each value inside the square root. This is just a tip. If you still doesnt know how to solve please tell me

OpenStudy (anonymous):

\[192=2^6\times 3\] \[243=3^5\] \[2^2\times 3^3\] this is a start

OpenStudy (anonymous):

I have no clue how to do this :[

OpenStudy (anonymous):

so for example \[\sqrt{192}=\sqrt{2^6}\times \sqrt{3}=2^3\times \sqrt{3}=8\times \sqrt{3}\]

OpenStudy (anonymous):

is that step clear?

OpenStudy (anonymous):

yes, but what do i do with the number infront?

OpenStudy (anonymous):

leave it there.

OpenStudy (anonymous):

you need to find a common value inside the square root so you can add the roots before. Is it clear?

OpenStudy (anonymous):

first term is \[-5\times 8\times \sqrt{3}=-40\sqrt{3}\]

OpenStudy (anonymous):

i haven't touched this stuff in like 4 months.

OpenStudy (anonymous):

second term is \[-2\times 3^2\times \sqrt{3}=-18\sqrt{3}\]

OpenStudy (anonymous):

third term is \[-8\times 2\times 3\times \sqrt{3}=-48\sqrt{3}\]

OpenStudy (anonymous):

then they are like terms so you can combine them

OpenStudy (anonymous):

you need to leave only a 3 inside each square root. 3 is the least common multiple number. so it must be something like: −5×(8×√3) , where (8×√3) = √192. Got it?

OpenStudy (anonymous):

yes thank you so much!

OpenStudy (anonymous):

@diego the point is not that 3 is the least "common multiple". the point is that when you write each in simplest radical form you end up with three terms that include \[\sqrt{3}\] your answer is of course right